# arXiv:2404.05610v1[cs.DC] 08 Apr 2024

### 08 Apr 2024 v1

@@ Abstract
We propose novel C++language bindings that cover all abstraction levels from low-level MPI calls to convenient STL-style bindings, where most parameters are inferred from a small subset of parameters, by bringing named parameters to C++. This enables rapid prototyping and fine-tuning runtime behavior and memory management. A flexible type system and additional safeness guarantees help to prevent programming errors.
By exploiting C++’s template-metaprogramming capabilities, this has (near) zero-overhead, as only required code paths are generated at compile time.
@@ I Introduction
The first version of the Message-Passing Interface (MPI) was proposed by the Message-Passing Interface Forum in 1994 [1] with the goal to standardize a portable, flexible, and efficient standard for message-passing. Today, it is the backbone of most HPC applications. While the majority of them is written in C++ [2], MPI’s syntax and semantics are designed around C and Fortran. While this allows for calling MPI from C++code, the semantics do not fit well with modern C++language features. This makes developing MPI application in C++unintuitive and error-prone [3].
@@
While all of them agree that introducing compatibility with STL containers, automatic data type deduction and an object oriented interface are key building blocks of such bindings, each new library chooses its own level of abstraction, requiring different amounts of boilerplate code. This may come at a performance penalty [8] and may introduce additional sources of errors.
To solve this, we introduce KaMPIng (Karlsruhe MPI next generation) [^1] https://github.com/kamping-site/kamping. Its main goal is to cover the complete range of abstraction levels over MPI calls, which makes it easy to use without introducing significant overhead compared to the C bindings. Parameters of MPI calls can either be provided directly by the user or are computed by KaMPIng. It further offers complete control over memory allocations.
It helps to reduce common sources of programming errors by employing compile-time error checking and prevents invalid memory access for nonblocking communication by introducing a ownership model. A flexible type system supports type-safety by generating type definitions at compile time and enables serialization when needed. Because all this is achieved using template meta-programming, only the code paths programmers would have to write themselves are generated, which makes these new bindings near zero-overhead.
We highlight the flexibility of our library by discussing a variety of different application benchmarks, including both simple examples and more complex applications scenarios. This includes applications in sorting, text processing, graph search and partitioning, and phylogenetic interference. For example, KaMPIng allows us to implement the Prefix Doubling algorithm for suffix array construction [9] using less than 200 lines of code and to entirely remove the custom MPI abstraction layer of RaxML-NG [10].
@@
  • Detailed overview over previous effort for designing C++MPI bindings.
  • A new approach to parameter handling allowing for a flexible computation of default parameters and fine-grained control over memory management for highly engineered distributed applications.
  • Safety for nonblocking MPI communication calls by design.
@@
  • (Near) zero-overhead compared to using plain MPI.
@@
The remainder of the paper is organized as follows. We begin by providing an overview of existing MPI (C++) language bindings in Section II. In Section III, we present KaMPIng’s core features like parameter handling, type deduction and serialization, handling of nonblocking communication, and our approach to ensure easy expandability. We then evaluate the applicability of KaMPIng in multiple different real-world application benchmarks in Section IV. Finally, we present first examples for general distributed computing algorithmic building blocks in Section V and conclude by discussing possibilities for future work in Section VI.
@@ II Related Work
Since the removal of C++bindings from the MPI standard, there has been continuous third-party effort on designing C++language bindings for MPI. The topic is also actively discussed in the MPI forum [^2] https://github.com/mpi-forum/mpi-issues/issues/288, where the following desired features where prominently mentioned:
@@
  • safety guarantees for nonblocking communication
@@
This was the first library to enable automatic data type inference and facilitate integration with the STL by also supporting std::vector for input and output, and not only raw pointers. Vectors are automatically resized to fit the received data, which prevents invalid memory accesses, but leads to hidden allocation. It supports custom data types by constructing appropriate data types if possible or resorting to serialization. To enable this, users must provide a serialization function compatible with Boost.Serialization, specifying all members explicitly. This requires the user to ensure that data type definitions and their serializations remain synchronized. The data types are managed by a global type pool which is queried for each communication call, which ensures proper resource cleanup. The adoption of Boost.MPI has been hindered by it’s tight coupling with other Boost libraries and the fact that it performs implicit serialization if data types are not directly supported by MPI [8]. Additionally, it is the only MPI library considered here that is not header-only. While this is usually not a problem with modern build tools, it becomes especially tedious when switching between MPI implementation, as it requires a separate build of Boost.MPI for each, because the MPI standard does not enforce ABI compatibility. Besides KaMPIng, it is the only library that supports mapping STL functors such as std::plus to the corresponding built-in MPI constant for reduction operations (which may enable optimization by the MPI implementation) and writing custom reduction operations using a simple lambda. If an MPI error occurs, an exception is thrown. Boost.MPI does not provide any bindings for MPI_Alltoallv. It is not actively maintained since 2008 and therefore only supports MPI-1.1 features. There exists a (non-official) rewrite of Boost.MPI, called boost-mpi3 [11], which aims to extend Boost.MPI to the feature set of MPI-3 and C++-17. While incorporating support for iterators, MPI one-sided communication and MPI shared memory, it does not substantially improve upon Boost.MPI’s design, closely mirroring MPI’s C interface.
@@
MPL introduces a powerful type system based on so called layouts to programmatically construct views over chunks of contiguous memory, which can be converted to MPI data types. While this allows for flexible communication in scientific computations, such as halo exchanges, it requires a lot of code for irregular communication patterns often found in discrete algorithms. MPL does not expose the underlying native MPI representation of constructs such as communicators, and types directly, which complicates iterative adaption of existing MPI code to MPL. MPL offers support for custom data types, which requires to define a matching layout, but has no serialization support. It currently does not support error handling.
Recently, MPL has been considered by the members of the newly formed MPI working group on language bindings as a starting point for new C++bindings [8]. While the authors highlight the simplified interface, which offers abstraction with default parameters using overloads, they also show that MPL incurs overheads for variable size collectives. This is due to the fact that it does not use the corresponding collective operations directly, but works with derived data types that are costly in some MPI implementations. For example, gatherv is implemented based on MPI_Alltoallw [8].
@@
Demiralp et al. recently introduced another modern C++interface, which we call RWTH-MPI in the following. They offer full STL support for send and receive buffers. For each communication call, they provide various overloads using different abstraction levels, which often allow the omission of send or receive counts, and RWTH-MPI employs additional communication to compute them. Automatic resizing of the receive buffer is supported in some cases and can be disabled. For custom types which support reflection based on the PFR library [12], RWTH-MPI constructs appropriate MPI data types automatically. A customization point for mapping C++types to type definitions for MPI is available, but does not provide automatic type management. Using types with dynamic run time sizes is not supported.
@@
There also exist MPI bindings for other high-level programming languages. The Python bindings provided by mpi4py [13, 14] may be considered as the most mature third-party bindings, as they are actively developed for over a decade. It enables transparent sending of Python objects using serialization. While this introduces additional overheads, using mpi4py in conjunction with NumPy arrays allows performance on par with a native C implementation [13]. They provide default values for certain parameters, but only when no additional communication is necessary for that. Finally, rsmpi [15] introduces idiomatic Rust bindings for MPI. While they make using MPI from Rust more ergonomic than using the C interface, the missing default parameters require the writing of much boilerplate code.
@@ III KaMPIng: Overview and Design
Similar to most previous bindings, KaMPIng represents MPI objects such as communicators, requests and statuses as classes and operations on them as member functions. Resource management is achieved using RAII [16, E.6][17] (Resource Acquisition Is Initialization) which is a commonly used C++idiom. Also, C++data types are mapped to MPI types at compile time, which prevents type matching errors. STL containers which allow access to the underlying contiguous memory are directly supported, i.e., every container which models the std::contiguous_range concept. Raw pointers are supported via std::span as proposed in the C++core guidelines [16, I.13]. One of KaMPIng’s distinct features is parameter handling. A common source of programming errors in MPI stems from the complexity of the interface of many communication calls. In particular, variable collective operations (suffixed with v) where the amount of data transferred between each processor pair varies, require a large number of parameters. The data to be sent or received is described in terms of a pointer to a memory region, the data type, the number of elements and its displacements. This makes MPI calls verbose and requires programmers to often consult the documentation for required parameters and ordering. While all parameters are necessary for full flexibility, there exist many use cases where only a small subset of explicitly provided parameters suffices and the remaining ones can be inferred from them. How we achieve this with near zero-overhead is discussed in Section III-A. In Section III-B, KaMPIng’s overall parameter handling concept is explained in more detail. If the number of elements to receive is already known, it may be desirable to resize containers appropriately, but for highly-tuned applications such hidden allocation may be unfavorable. We therefore propose a flexible allocation control in Section III-C, which is configurable at compile time.
@@
Non-blocking communication in MPI introduces additional sources of errors, as the user has to manually wait for the completion of operations and has to take care to not perform invalid memory accesses before a operation has finished. To address this, we propose memory-safe abstractions which prevent illegal memory accesses through the use of C++’s ownership model and move semantics in Section III-E.
@@ III-A Computation of Default Parameters
As discussed previously, MPI calls often allow for computing useful default values for an operation based on only a small subset of parameters. As an example, consider the case, where we want to perform an MPI_Allgatherv, where each rank initially holds an std::vector of varying size and we want to concatenate them to a global vector on each rank. The send count and data type can be directly inferred from the vector’s size and value_type (see Section III-D for more details on automatic type deduction). Receive counts and displacements can be computed by an MPI_Allgather of all send counts followed by an exclusive prefix sum over them (see Figure 1). While this is a common pattern, none of the other C++bindings allows to avoid all this boilerplate code. Boost.MPI offers various overloaded functions which allow the user to omit explicit displacements, but the counts have to be communicated. RWTH-MPI does provide an overload which gathers the counts internally, but it only works with MPI_IN_PLACE, which requires the send data to be already provided at the correct position on each rank. To achieve this the user has to manually exchange count information upfront  [^3] Full example code can be found at https://github.com/kamping-site/kamping-examples/tree/main/include/vector_allgather/. This leaves us with a situation where the usability of C++bindings depends on whether the implementers had this particular use case in mind and provided a default option for it.
To address this problem, we choose an alternative approach inspired by named parameters, where parameters passed to a function can be named at the caller site and passed in arbitrary order (as known from languages like Python). Internally, named parameters are realized as factory functions [18] which construct lightweight parameter objects encapsulating the parameter type (i.e., send buffer, send counts, $\dots$) and the corresponding data. This allows us to check for the presence of each parameter at compile time and to compute default values only if the respective parameter is omitted, without resorting to many overloads exploring the complete combinatorial explosion of parameters. To avoid runtime overhead, we rely on template meta-programming to only generate the code paths required for computing missing parameters at compile time. Gathering a vector in KaMPIng then becomes a one-liner as shown in Figure 1. The implementations using other bindings are more verbose (see Footnote 6).
@@
//allgathering a vector using MPI
int size;
@@
int n = static_cast<int>(v.size());
@@
MPI_Allgather(&n, 1, MPI_INT, rc.data(), 1, MPI_INT, comm);
@@
//allgathering a vector using __MPIng
@@
Figure 1: Allgathering of a vector using MPI vs. KaMPIng.

### 23 Aug 2024 v2

# arXiv:2404.05610v2[cs.DC] 23 Aug 2024

@@ Abstract
We propose novel C++ language bindings that cover all abstraction levels from low-level MPI calls to convenient STL-style bindings, where most parameters are inferred from a small subset of parameters, by bringing named parameters to C++. This enables rapid prototyping and fine-tuning runtime behavior and memory management. A flexible type system and additional safety guarantees help to prevent programming errors.
By exploiting C++’s template metaprogramming capabilities, this has (near) zero overhead, as only required code paths are generated at compile time.
@@ I Introduction
The first version of the Message-Passing Interface (MPI) was proposed by the Message-Passing Interface Forum in 1994 [1] with the goal of standardizing a portable, flexible, and efficient standard for message-passing. Today, it is the backbone of most HPC applications. While the majority of them are written in C++ [2], MPI’s syntax and semantics are designed around C and Fortran. While this allows for calling MPI from C++ code, the semantics do not fit well with modern C++ language features. This makes developing MPI applications in C++ unintuitive and error-prone [3].
@@
Previous bindings focussed on making MPI’s C interface compatible with C++ features such as templates, STL containers and object orientation, and providing some sensible defaults, but this resulted in each library choosing its own level of abstraction: They either provide a high-level interface, while often sacrificing performance [8], or stay close to MPI’s C interface, still requiring large amounts of boilerplate code. Additionally, these previous libraries do not work well with move semantics and common practices proposed by the C++ core guidelines, e.g., returning results by value instead of using C-style in/out parameters [9, F.20]. Despite the previous efforts, designing a better C++ MPI interface is an open problem actively discussed in the MPI forum [10].
Therefore, we propose the C++ MPI binding library KaMPIng (Karlsruhe MPI next generation) [^1] https://github.com/kamping-site/kamping, which we recently announced briefly at SPAA [11]. Its main goal is to cover the complete range of abstraction levels over MPI calls as shown in Figure 1: KaMPIng enables rapid prototyping, relying on its sensible defaults, as well as engineering highly-tuned distributed code by making use of the library’s flexible interface. By using C++ template metaprogramming techniques this comes without introducing significant overhead compared to MPI’s plain C bindings.
Parameters of MPI calls can either be provided directly by the user or are computed by KaMPIng. It further offers complete control over memory allocations.
It helps to reduce common sources of programming errors by employing compile-time error checking and prevents invalid memory access for non-blocking communication by introducing an ownership model. A flexible type system supports type safety by generating type definitions at compile time and enables serialization when needed. Because all this is achieved using template metaprogramming, only the code paths programmers would have to write themselves are instantiated, while the rest can be eliminated at compile-time, which makes these new bindings near zero-overhead.
Currently, KaMPIng supports the most commonly used MPI features [2], namely (non-blocking) point-to-point communication and collective communication. Moreover, its core architecture is designed with the rest of the MPI standard in mind, facilitating a straightforward implementation of it in the future.
We highlight the flexibility of our library by discussing a variety of different application benchmarks, including both simple examples and more complex application scenarios. This includes applications in sorting, text processing, graph search and partitioning, and phylogenetic interference. For example, KaMPIng allows us to implement the Prefix Doubling algorithm for suffix array construction [12] using less than 200 lines of code and to entirely remove the custom MPI abstraction layer of RaxML-NG [13].
std::vector<double> v = {...};
// KaMPIng allows concise code
// with sensible defaults ... (1)
auto v_global = comm.allgatherv(send_buf(v));
// ... or detailed tuning of each parameter (2)
std::vector<int> rc;
auto [v_global, rdispls] = comm.allgatherv(
send_buf(v), //(3)
recv_counts_out<resize_to_fit/*(6)*/>(std::move(rc)), //(4)
recv_displs_out() //(5)
Figure 1: KaMPIng offers a high-level easy-to-use interface (1) and full control over each parameter (2). Data types and buffer sizes can be automatically inferred (3). Arguments allow passing by reference or by value and transferring data ownership via move semantics (4). Out-parameters allow controlling which default-computed parameters are returned to the user (5). Resize policies allow controlling memory allocation (6).
@@
  • A new approach to parameter handling, allowing for a flexible computation of default parameters and fine-grained control over memory management for highly engineered distributed applications.
  • Safety for non-blocking MPI communication calls by design.
@@
  • (Near) zero overhead compared to using plain MPI.
@@
The remainder of the paper is organized as follows. We begin by providing an overview of existing MPI (C++) language bindings in Section II. In Section III, we present KaMPIng’s core features like parameter handling, type deduction and serialization, handling of non-blocking communication, and our approach to ensure easy expandability. We then evaluate the applicability of KaMPIng in multiple different real-world application benchmarks in Section IV. Finally, we present first examples for general distributed computing algorithmic building blocks in Section V and conclude by discussing possibilities for future work in Section VI.
@@ II Related Work
Since the removal of C++ bindings from the MPI standard, there has been continuous third-party effort on designing C++ language bindings for MPI. The topic is also actively discussed in the MPI forum [10], where the following desired features were prominently mentioned:
@@
  • safety guarantees for non-blocking communication
@@
This was the first library to enable automatic data type inference and facilitate integration with the STL by also supporting std::vector for input and output, and not only raw pointers. Vectors are automatically resized to fit the received data, which prevents invalid memory accesses, but leads to hidden allocation. It supports custom data types by constructing appropriate data types if possible or resorting to serialization. To enable this, users must provide a serialization function compatible with Boost.Serialization, specifying all members explicitly. This requires the user to ensure that data type definitions and their serializations remain synchronized. The data types are managed by a global type pool which is queried for each communication call, which ensures proper resource cleanup. The adoption of Boost.MPI has been hindered by its tight coupling with other Boost libraries and the fact that it performs implicit serialization if data types are not directly supported by MPI [8]. Additionally, it is the only MPI library considered here that is not header-only. While this is usually not a problem with modern build tools, it becomes especially tedious when switching between MPI implementation, as it requires a separate build of Boost.MPI for each, because the MPI standard does not enforce ABI compatibility. Besides KaMPIng, it is the only library that supports mapping STL functors such as std::plus to the corresponding built-in MPI constant for reduction operations (which may enable optimization by the MPI implementation) and writing custom reduction operations using a simple lambda. If an MPI error occurs, an exception is thrown. Boost.MPI does not provide any bindings for MPI_Alltoallv. It has not been actively maintained since 2008 and therefore only supports MPI-1.1 features. There also exists boost-mpi3 [14], a rewrite of Boost.MPI developed independently of the Boost project but following its design principles. It aims to extend Boost.MPI to the feature set of MPI-3 and C++-17. While incorporating support for iterators, MPI one-sided communication and MPI shared memory, it does not substantially improve upon Boost.MPI’s design, but closely mirrors MPI’s C interface.
@@
MPL introduces a powerful type system based on so-called layouts to programmatically construct views over chunks of contiguous memory, which can be converted to MPI data types. While this allows for flexible communication in scientific computations, such as halo exchanges, it requires a lot of code for irregular communication patterns often found in discrete algorithms. MPL does not expose the underlying native MPI representation of constructs such as communicators, and types directly, which complicates the iterative adaption of existing MPI code to MPL. MPL offers support for custom data types, which requires defining a matching layout but has no serialization support. It currently does not support error handling.
Recently, MPL has been considered by the members of the newly formed MPI working group on language bindings as a starting point for new C++ bindings [8]. While the authors highlight the simplified interface, which offers abstraction with default parameters using overloads, they also show that MPL incurs overheads for variable-size collectives. This is because it does not use the corresponding collective operations directly by passing appropriate counts/displacements, but constructs custom derived data types with displacements. Therefore, operations like gatherv call MPI_Alltoallw internally, which is costly in some MPI implementations and limits scalability [8].
@@
Demiralp et al. recently introduced another modern C++ interface, which we call RWTH-MPI in the following. They offer full STL support for send and receive buffers. For each communication call, they provide various overloads using different abstraction levels, which often allow the omission of send or receive counts, and RWTH-MPI employs additional communication to compute them. Automatic resizing of the receive buffer is supported in some cases and can be disabled. For custom types that support reflection based on the PFR library [15], RWTH-MPI constructs appropriate MPI data types automatically. A customization point for mapping C++ types to type definitions for MPI is available but does not provide automatic type management. Using types with dynamic run time sizes is not supported.
@@
There also exist MPI bindings for other high-level programming languages. The Python bindings provided by mpi4py [16, 17] may be considered the most mature third-party bindings, as they have been actively developed for over a decade. It enables the transparent sending of Python objects using serialization. While this introduces additional overheads, using mpi4py in conjunction with NumPy arrays allows performance on par with a native C implementation [16]. They provide default values for certain parameters, but only when no additional communication is necessary for that. Finally, rsmpi [18] introduces idiomatic Rust bindings for MPI. While they make using MPI from Rust more ergonomic than using the C interface, the missing default parameters require the writing of much boilerplate code.
@@ III KaMPIng: Overview and Design
Realizing a C++ interface that is both easy to use and highly flexible, requires building a library from scratch, improving upon the design of its precursors. Similar to most previous bindings, KaMPIng represents MPI objects such as communicators, requests and statuses as classes and operations on them as member functions. Automatic creation/destruction of MPI objects is achieved using RAII [9, E.6][19] (Resource Acquisition Is Initialization) which is a commonly used C++ idiom. Also, C++ data types are mapped to MPI types at compile time, which prevents type-matching errors. STL containers that allow access to the underlying contiguous memory are directly supported, i.e., every container that models the std::contiguous_range concept. Raw pointers are supported via std::span as proposed in the C++ core guidelines [9, I.13]. One of KaMPIng’s distinct features is parameter handling. A common source of programming errors in MPI stems from the complexity of the interface of many communication calls. In particular, variable collective operations (suffixed with v) where the amount of data transferred between each processor pair varies, require a large number of parameters. The data to be sent or received is described in terms of a pointer to a memory region, the data type, the number of elements and its displacements. This makes MPI calls verbose and requires programmers to often consult the documentation for required parameters and ordering. While all parameters are necessary for full flexibility, there exist many use cases where only a small subset of explicitly provided parameters suffices and the remaining ones can be inferred from them. How we achieve this with near zero overhead is discussed in Section III-A. In Section III-B, KaMPIng’s overall parameter handling concept is explained in more detail. If the number of elements to receive is already known, it may be desirable to resize containers appropriately, but for highly-tuned applications such hidden allocation may be unfavorable. We therefore propose a flexible allocation control in Section III-C, which is configurable at compile time.
@@
Non-blocking communication in MPI introduces additional sources of errors, as the user has to manually wait for the completion of operations and has to take care of not performing invalid memory accesses before an operation has finished. To address this, we propose memory-safe abstractions that prevent illegal memory accesses through the use of C++’s ownership model and move semantics in Section III-E.
@@
In Section III-G we describe how KaMPIng prevents common programming errors by providing many compile- and runtime assertions and configurable error handling, and by introducing a simplified syntax for in-place operations. Finally, Section III-H details how KaMPIng is implemented internally.
@@ III-A Computation of Default Parameters
int size, rank;
@@
MPI_Comm_rank(comm, &rank);
@@
rc[rank] = v.size();
//exchange counts
MPI_Allgather(MPI_IN_PLACE, 0, MPI_DATATYPE_NULL,
rc.data(), 1, MPI_INT, comm);
//compute displacements
@@
Figure 2: Allgathering of a vector using MPI.
As discussed previously, MPI calls often allow for computing useful default values for an operation based on only a small subset of parameters. As an example, consider the case, where we want to perform an MPI_Allgatherv, where each rank initially holds an std::vector of varying size and we want to concatenate them to a global vector on each rank. The send count and data type can be directly inferred from the vector’s size and value_type (see Section III-D for more details on automatic type deduction). Receive counts and displacements can be computed by an MPI_Allgather of all send counts followed by an exclusive prefix sum over them (see Figure 2). While this is a common pattern, none of the other C++ bindings allows avoiding all this boilerplate code. Boost.MPI offers various overloaded functions that allow the user to omit explicit displacements, but the counts have to be communicated. RWTH-MPI does provide an overload that gathers the counts internally, but it only works with MPI_IN_PLACE, which requires the send data to be already provided at the correct position on each rank. To achieve this the user has to manually exchange count information upfront [^2] Full example code can be found at https://github.com/kamping-site/kamping-examples/tree/main/include/vector_allgather/. This leaves us with a situation where the usability of C++ bindings depends on whether the implementers had this particular use case in mind and provided a default option for it.
To address this problem, we choose an alternative approach inspired by named parameters, where parameters passed to a function can be named at the caller site and passed in arbitrary order (as known from languages like Python). Internally, named parameters are realized as factory functions [20] which construct lightweight parameter objects encapsulating the parameter type (i.e., send buffer, send counts, $\dots$) and the corresponding data. This allows us to check for the presence of each parameter at compile time and to compute default values only if the respective parameter is omitted, without resorting to many overloads exploring the complete combinatorial explosion of parameters. To avoid runtime overhead, we rely on template metaprogramming to only generate the code paths required for computing missing parameters at compile time. Gathering a vector in KaMPIng then becomes a one-liner as shown in Figure 1. The implementations using other bindings are more verbose (see Footnote 7).
The flexibility of the named parameter interface allows users to iteratively adapt their existing code to KaMPIng, as shown in Figure 3.
// Version 1: using KaMPIng’s interface
std::vector<int> rc(comm.size()), rd(comm.size());
rc[comm.rank()] = v.size();
comm.allgather(send_recv_buf(rc));
std::exclusive_scan(rc.begin(), rc.end(), rd.begin(), 0);
std::vector<T> v_glob;
comm.allgatherv(send_buf(v), recv_buf(v_glob),
recv_counts(rc), recv_displs(rd));
// Version 2: displacements are computed implicitly
std::vector<int> rc(comm.size());
rc[comm.rank()] = v.size();
comm.allgather(send_recv_buf(rc));
std::vector<T> v_glob;
comm.allgatherv(send_buf(v),
recv_buf<resize_to_fit>(v_glob), recv_counts(rc));
// Version 3: counts are automatically exchanged
// and result is returned by value
@@
Figure 3: The vector allgather MPI example from Figure 2 can be gradually migrated to use more of KaMPIng’s features.